OpenStudy (anonymous):

cosx cos (pi/7)+sinx sin (pi/7)= -sqrt2/2 x can equal:

OpenStudy (anonymous):

\[5\pi/4+\pi/7+2n \pi\]

OpenStudy (anonymous):

\[7\pi/4+\pi/7+2n \pi\]

OpenStudy (anonymous):

wat is rhs?

OpenStudy (anonymous):

\[3\pi/4+\pi/7+2n \pi\]

OpenStudy (anonymous):

\[\cos x \cos (\pi/7)+\sin xsin(\pi/7)=\cos (x-\pi/7)=-\sqrt{2}/2\]

OpenStudy (anonymous):

1 nd 3

OpenStudy (anonymous):

or \[\pi/4+\pi/7+2n \pi\]

OpenStudy (anonymous):

corrected

OpenStudy (anonymous):

which of the 4 equations do you think is correct? there can be more than 1

OpenStudy (anonymous):

1st nd 3rd

OpenStudy (anonymous):

thank youu :)